OpenStudy (anonymous):

Find d2y/dx^2, in terms of x and y for the equation y^2=x^3

OpenStudy (anonymous):

first you need y'

OpenStudy (anonymous):

\[2yy'=3x^2\]

OpenStudy (anonymous):

or \[y'=\frac{3x^2}{2y}\] now again, this time with quotient rule

OpenStudy (anonymous):

constants getting on my nerves. lets write it as \[\frac{3}{2}\frac{x^2}{y}\] and take the derivative of second part. you get \[y''=\frac{2}{3}\frac{2xy-x^2y'}{y^2}\] so far so good?

OpenStudy (anonymous):

now replace \[y'\] by \[\frac{3x^2}{2y}\] and do some annoying algebra to get your answer

OpenStudy (anonymous):

\[\frac{2}{3}\frac{2xy-x^2\times \frac{3x^2}{2y}}{y^2}\] multiply top and bottom by 2y to clear the fraction get \[\frac{2}{3}\frac{4xy^2-3x^4}{2y^3}\] check my algebra because i am typing not writing.

OpenStudy (anonymous):

now don't forget that \[y^2=x^3\] so you can replace all y squares you see by x cubed and simplify more

OpenStudy (anonymous):

yo satellite please help me when you're done here

OpenStudy (anonymous):

guess i am done